# Preference-based Reinforcement Learning beyond Pairwise Comparisons: Benefits of Multiple Options

### v1

@@ D.2 Proofs of Lemmas for Theorem 1
$\sum_{j=1}^{|S_{s}|}\nabla^{2}\ell^{(j)}_{s}(\bm{\theta}^{\star})$ $=\sum_{j=1}^{|S_{s}|}\sum_{k=j}^{|S_{s}|}\sum_{k^{\prime}=j}^{|S_{s}|}\frac{\exp\left(\left(\phi(x_{s},\sigma_{sk})+\phi(x_{s},\sigma_{sk^{\prime}})\right)^{\top}\bm{\theta}^{\star}\right)}{2\left(\sum_{k^{\prime}=j}^{|S_{s}|}\exp\left(\phi(x_{s},\sigma_{sk^{\prime}})^{\top}\bm{\theta}^{\star}\right)\right)^{2}}\cdot z_{skk^{\prime}}z_{skk^{\prime}}^{\top}$
$=\frac{1}{2}\sum_{j=1}^{|S_{s}|}\sum_{k=j}^{|S_{s}|}\sum_{k^{\prime}=j}^{|S_{s}|}P_{s}(\sigma_{sk}|S_{s}^{(j)};\bm{\theta}^{\star})P_{s}(\sigma_{sk^{\prime}}|S_{s}^{(j)};\bm{\theta}^{\star})z_{skk^{\prime}}z_{skk^{\prime}}^{\top}$
$=\frac{1}{2}\sum_{j=1}^{|S_{s}|}\mathbb{E}_{(a,a^{\prime})\sim P_{s}^{(j)}\times P_{s}^{(j)}}\left[\big(\phi(x_{s},a)-\phi(x_{s},a^{\prime})\big)\big(\phi(x_{s},a)-\phi(x_{s},a^{\prime})\big)^{\top}\right],$ (D.10)
where $z_{skk^{\prime}}=\phi(x_{s},\sigma_{sk})-\phi(x_{s},\sigma_{sk^{\prime}})$. Let the action $\bar{a}_{s}\in S_{s}$ be ranked at position $\bar{k}_{s}$ in the ranking $\sigma_{s}$. That is,
$\sigma_{s}=(\underbrace{\sigma_{s1},\dots\sigma_{s\bar{k}_{s}-1}}_{\bar{k}_{s}-1\text{ actions}},\bar{a}_{s},\sigma_{s\bar{k}_{s}+1},\dots\sigma_{s|S_{s}|-1}).$
Note that $\bar{a}_{s}\in S_{s}^{(j)}$ for $j\leq\bar{k}_{s}$. We also note that $P_{s}^{(j)}$ is measurable with respect to the filtration $\mathcal{F}^{\prime}_{s-1,j-1}=\bm{\sigma}\left(S_{1},\sigma_{11},\sigma_{12},\dots,S_{s},\sigma_{s1},\dots\sigma_{sj-1}\right)$ . Then, by plugging Equation (D.10) into Equation (D.9) and applying the covariance matrix concentration result (Corollary F.1), since $\lambda=\Omega(d\log(KT/\delta))$, we have, with probability at least $1-\delta$,
$H_{t}$ $\succeq\frac{1}{2e}\sum_{s\in[t-1]\setminus{\mathcal{T}}^{w}}\sum_{j=1}^{|S_{s}|}\mathbb{E}_{(a,a^{\prime})\sim P_{s}^{(j)}\times P_{s}^{(j)}}\left[\big(\phi(x_{s},a)-\phi(x_{s},a^{\prime})\big)\big(\phi(x_{s},a)-\phi(x_{s},a^{\prime})\big)^{\top}\right]+\lambda\mathbf{I}_{d}$
$\succeq\frac{1}{2e}\sum_{s\in[t-1]\setminus{\mathcal{T}}^{w}}\sum_{j=1}^{\bar{k}_{s}}\mathbb{E}_{(a,a^{\prime})\sim P_{s}^{(j)}\times P_{s}^{(j)}}\left[\big(\phi(x_{s},a)-\phi(x_{s},a^{\prime})\big)\big(\phi(x_{s},a)-\phi(x_{s},a^{\prime})\big)^{\top}\right]+\lambda\mathbf{I}_{d}$ ($\bar{k}_{s}\leq|S_{s}|$)
$\succeq\frac{3}{10e}\left(\sum_{s\in[t-1]\setminus{\mathcal{T}}^{w}}\sum_{j=1}^{\bar{k}_{s}}\big(\phi(x_{s},\sigma_{sj})-\phi(x_{s},\bar{a}_{s})\big)\big(\phi(x_{s},\sigma_{sj})-\phi(x_{s},\bar{a}_{s})\big)^{\top}+\lambda\mathbf{I}_{d}\right)$ (Corollary F.1, $\bar{a}_{s}\in S_{s}^{(j)}$ for $j\leq\bar{k}_{s}$ )
$=\frac{3K}{10e}\Bigg(\sum_{s\in[t-1]\setminus{\mathcal{T}}^{w}}\underbrace{\frac{1}{K}\sum_{j=1}^{\bar{k}_{s}}\big(\phi(x_{s},\sigma_{sj})-\phi(x_{s},\bar{a}_{s})\big)\big(\phi(x_{s},\sigma_{sj})-\phi(x_{s},\bar{a}_{s})\big)^{\top}}_{=:X(\sigma_{s})}+\lambda\mathbf{I}_{d}\Bigg).$
Here, $\{X(\sigma_{s})\}_{s\in[t-1]\setminus{\mathcal{T}}^{w}}$ is a sequence of positive semi-definite (PSD) random matrices, where each matrix $X(\sigma_{s})$ depends on the sampled ranking $\sigma_{s}$, and satisfies $\lambda_{\max}\left(X(\sigma_{s})\right)\leq 1$.
@@
Note that the ranking $\sigma_{s}$ is drawn from the PL distribution $\mathbb{P}(\cdot\mid x_{s},S_{s};\bm{\theta}^{\star})$, which is measurable with respect to the filtration $\mathcal{F}_{s-1}=\bm{\sigma}(S_{1},\sigma_{1},\dots,S_{s})$. Furthermore, $X(\sigma_{s})$ is measurable with respect to $\bm{\sigma}\left(\mathcal{F}_{s-1},\sigma_{s}\right)$. Then, by applying the concentration lemma for PSD matrices (Lemma F.4) two times, with probability at least $1-2\delta$, we get
@@
$H_{t}$ $\succeq\frac{3K}{10e}\left(\sum_{s\in[t-1]\setminus{\mathcal{T}}^{w}}X(\sigma_{s})+\lambda\mathbf{I}_{d}\right)$
$\succeq\frac{K}{10e}\left(\sum_{s\in[t-1]\setminus{\mathcal{T}}^{w}}\mathbb{E}_{\sigma\sim\mathbb{P}(\cdot|x_{s},S_{s};\bm{\theta}^{\star})}\left[X(\sigma)\right]+\lambda\mathbf{I}_{d}\right)$ (Lemma F.4)
$\succeq\frac{3K}{50e}\left(\sum_{s\in[t-1]\setminus{\mathcal{T}}^{w}}X(\tilde{\sigma}_{s})+\lambda\mathbf{I}_{d}\right),$ (Lemma F.4)
where $\tilde{\sigma}_{s}$ denotes an arbitrary ranking in which $\bar{a}_{s}$ is placed last. For example, $\tilde{\sigma}_{s}=(\sigma_{s1},\dots,\sigma_{s\bar{k}_{s}-1},\sigma_{s\bar{k}_{s}+1},\sigma_{s|S_{s}|-1},\bar{a}_{s})$. Note that $\tilde{\sigma}_{s}$ is a possible virtual ranking feedback for the assortment $S_{s}$, whereas $\sigma_{s}$ denotes the actual ranking feedback observed at round $s$. Hence, since $\bar{a}_{s}$ occupies the final position in the virtual sequence $\tilde{\sigma}_{s}$, it follows that:
$H_{t}$ $\succeq\frac{3K}{50e}\left(\sum_{s\in[t-1]\setminus{\mathcal{T}}^{w}}X(\tilde{\sigma}_{s})+\lambda\mathbf{I}_{d}\right)$
$=\frac{3}{50e}\left(\sum_{s\in[t-1]\setminus{\mathcal{T}}^{w}}\sum_{j=1}^{|S_{s}|}\big(\phi(x_{s},\tilde{\sigma}_{sj})-\phi(x_{s},\bar{a}_{s})\big)\big(\phi(x_{s},\tilde{\sigma}_{sj})-\phi(x_{s},\bar{a}_{s})\big)^{\top}+\lambda\mathbf{I}_{d}\right)$
@@
$=\frac{3}{50e}\left(\sum_{s\in[t-1]\setminus{\mathcal{T}}^{w}}\sum_{a\in S_{s}}\big(\phi(x_{s},a)-\phi(x_{s},\bar{a}_{s})\big)\big(\phi(x_{s},a)-\phi(x_{s},\bar{a}_{s})\big)^{\top}+\lambda\mathbf{I}_{d}\right)$
$=\frac{3}{50e}\Lambda_{t}\succeq\frac{1}{50}\Lambda_{t}.$ (Def. of $\Lambda_{t}$, Eqn. (D.2))
By substituting $\delta\leftarrow\frac{\delta}{3}$, we conclude the proof of Lemma D.2. ∎
D.2.2 Proof of Lemma D.3
Proof of Lemma D.3.
By the definition of $\Lambda_{t}$, we have
$\operatorname{det}\left(\Lambda_{t+1}\right)$ $=\operatorname{det}\left(\Lambda_{t}+\sum_{a\in S_{t}}z_{ta}z_{ta}^{\top}\right)$
$\geq\operatorname{det}\left(\Lambda_{t}\right)\left(1+\sum_{a\in S_{t}}\|z_{ta}\|_{\Lambda_{t}^{-1}}^{2}\right)$
$\geq\operatorname{det}\left(\lambda\mathbf{I}_{d}\right)\prod_{s=1}^{t}\left(1+\sum_{a\in S_{s}}\|z_{sa}\|_{\Lambda_{s}^{-1}}^{2}\right)$
$\geq\operatorname{det}\left(\lambda\mathbf{I}_{d}\right)\prod_{s=1}^{t}\left(1+\min\left\{1,\sum_{a\in S_{s}}\|z_{sa}\|_{\Lambda_{s}^{-1}}^{2}\right\}\right).$ (D.11)
Then, using the fact that $a\leq 2\log(1+a)$ for any $a\in[0,1]$, we get
$\sum_{t=1}^{T}\min\left\{1,\sum_{a\in S_{t}}\|z_{ta}\|_{\Lambda_{t}^{-1}}^{2}\right\}$ $\leq 2\sum_{t=1}^{T}\log\left(1+\min\left\{1,\sum_{a\in S_{t}}\|z_{ta}\|_{\Lambda_{t}^{-1}}^{2}\right\}\right)$
$\leq 2\log\left(\frac{\operatorname{det}\left(\Lambda_{T+1}\right)}{\operatorname{det}\left(\lambda\mathbf{I}_{d}\right)}\right)$ (Eqn. (D.11))
$\leq 2d\log\left(1+\frac{X^{2}KT}{d\lambda}\right),$
@@
$\operatorname{det}\left(\Lambda_{T+1}\right)$ $\leq\left(\frac{\lambda_{1}+\cdots+\lambda_{d}}{d}\right)^{d}$ ($\lambda_{1},\cdots,\lambda_{d}$ are eigenvalues of $\Lambda_{T+1}$, AM-GM ineq.)
$=\left(\frac{\operatorname{trace}(\Lambda_{T+1})}{d}\right)^{d}$
$=\left(\frac{\lambda d+\sum_{t=1}^{T}\sum_{a\in S_{t}}\|z_{ta}\|_{2}^{2}}{d}\right)^{d}\leq\left(\lambda+\frac{X^{2}KT}{d}\right)^{d}.$
This concludes the proof of Lemma D.3. ∎
D.2.3 Proof of Lemma D.4
Proof of Lemma D.4.
Let $W_{t}:=\lambda\mathbf{I}_{d}+\sum_{s\in{\mathcal{T}}_{0},s<t}\sum_{a\in S_{s}}z_{sa}z_{sa}^{\top}+\lambda\mathbf{I}_{d}$. Then, we have
$\left(\lambda+\frac{X^{2}|{\mathcal{T}}_{0}|K}{d}\right)^{d}$ $\geq\left(\frac{\lambda d+\sum_{t\in{\mathcal{T}}_{0}}\sum_{a\in S_{t}}\|z_{ta}\|_{2}^{2}}{d}\right)^{d}$
$=\left(\frac{\operatorname{trace}(W_{T+1})}{d}\right)^{d}$
$\geq\operatorname{det}(W_{T+1})$ (AM-GM ineq.)
$=\operatorname{det}(\lambda\mathbf{I}_{d})\prod_{t\in{\mathcal{T}}_{0}}\left(1+\sum_{a\in S_{t}}\|z_{ta}\|_{W_{t}^{-1}}^{2}\right)$ (update equality for det.)
$\geq\operatorname{det}(\lambda\mathbf{I}_{d})\prod_{t\in{\mathcal{T}}_{0}}\left(1+\sum_{a\in S_{t}}\|z_{ta}\|_{\Lambda_{t}^{-1}}^{2}\right)$ ($W_{t}\preceq\Lambda_{t}$)
$\geq\lambda^{d}(1+L)^{|{\mathcal{T}}_{0}|}.$ ($\sum_{a\in S_{t}}\|z_{ta}\|_{\Lambda_{t}^{-1}}^{2}\geq L$ for $t\in{\mathcal{T}}_{0}$)
Hence, we get
$|{\mathcal{T}}_{0}|$ $\leq\frac{d}{\log(1+L)}\log\left(1+\frac{X^{2}|{\mathcal{T}}_{0}|K}{d\lambda}\right)$ (D.12)
$=\frac{d}{\log(1+L)}\Bigg(\log\left(\frac{|{\mathcal{T}}_{0}|}{2d/\log(1+L)}\right)+\log\left(\frac{2d}{\log(1+L)}\left(\frac{1}{|{\mathcal{T}}_{0}|}+\frac{X^{2}K}{d\lambda}\right)\right)\Bigg)$
$\leq\frac{|{\mathcal{T}}_{0}|}{2}+\frac{d}{\log(1+L)}\log\left(\frac{2d}{e\log(1+L)}\left(\frac{1}{|{\mathcal{T}}_{0}|}+\frac{X^{2}K}{d\lambda}\right)\right),$
which implies that
$|{\mathcal{T}}_{0}|\leq\frac{2d}{\log(1+L)}\log\left(\frac{2d}{e\log(1+L)}\left(\frac{1}{|{\mathcal{T}}_{0}|}+\frac{X^{2}K}{d\lambda}\right)\right).$ (D.13)
Now, we fix $c>0$ and consider two cases:
  • Case 1: $|{\mathcal{T}}_{0}|<cd$
    In this case, from Equation (D.12), we have $|{\mathcal{T}}_{0}|\leq\frac{d}{\log(1+L)}\log\left(1+\frac{X^{2}cK}{\lambda}\right)$.
  • Case 2: $|{\mathcal{T}}_{0}|\geq cd$
    In this case, from Equation (D.13), we have $|{\mathcal{T}}_{0}|\leq\frac{2d}{\log(1+L)}\log\left(\frac{2}{e\log(1+L)}\left(\frac{1}{c}+\frac{X^{2}K}{\lambda}\right)\right)$.
By setting $c=\frac{2}{e\log(1+L)}$, we obtain
$|{\mathcal{T}}_{0}|\leq\frac{2d}{\log(1+L)}\log\left(1+\frac{X^{2}K}{\log(1+L)\lambda}\right),$
which concludes the proof of Lemma D.4. ∎
D.2.4 Proof of Lemma D.5
Proof of Lemma D.5.
For simplicity, let $\widetilde{{\mathcal{T}}}^{w}_{t}=\{s\in[t-1]\mid s\in{\mathcal{T}}^{w}\cap({\mathcal{T}}_{0})^{c}\}$. Clearly, $\widetilde{{\mathcal{T}}}^{w}_{T+1}={\mathcal{T}}^{w}\cap({\mathcal{T}}_{0})^{c}$. Recall that by the definition of $H_{t}$, we have
@@
$\succeq\frac{\kappa}{2K^{2}}\sum_{s=1}^{t-1}\sum_{k=j}^{|S_{s}|}\sum_{k^{\prime}=j}^{|S_{s}|}z_{skk^{\prime}}z_{skk^{\prime}}^{\top}+\lambda\mathbf{I}_{d}$ ($\kappa=e^{-4B}$)
$\succeq\frac{\kappa}{2K^{2}}\sum_{s=1}^{t-1}\sum_{a\in S_{s}}\big(\phi(x_{s},a)-\phi(x_{s},\bar{a}_{s})\big)\big(\phi(x_{s},a)-\phi(x_{s},\bar{a}_{s})\big)^{\top}+\lambda\mathbf{I}_{d}$ ($\bar{a}_{s}\in S_{s}$ by Eqn. (9))
$\succeq\frac{\kappa}{2K^{2}}\,\underbrace{\left(\sum_{s\in\widetilde{{\mathcal{T}}}^{w}_{t}}\sum_{a\in S_{s}}\big(\phi(x_{s},a)-\phi(x_{s},\bar{a}_{s})\big)\big(\phi(x_{s},a)-\phi(x_{s},\bar{a}_{s})\big)^{\top}+\lambda\mathbf{I}_{d}\right)}_{=:\Lambda^{w}_{t}},$ (D.14)
where $z_{skk^{\prime}}=\phi(x_{s},\sigma_{sk})-\phi(x_{s},\sigma_{sk^{\prime}})$.
Let $\tilde{a}_{t}=\mathop{\mathrm{argmax}}_{a\in\mathcal{A}}\|\phi(x_{t},a)-\phi(x_{t},\bar{a}_{t})\|_{H_{t}^{-1}}$. Then, we get
$\sum_{t\in\widetilde{{\mathcal{T}}}^{w}_{T+1}}\max_{a\in\mathcal{A}}$ $\|\phi(x_{t},a)-\phi(x_{t},\bar{a}_{t})\|^{2}_{H_{t}^{-1}}$
$\leq\sum_{t\in\widetilde{{\mathcal{T}}}^{w}_{T+1}}\|\phi(x_{t},\tilde{a}_{t})-\phi(x_{t},\bar{a}_{t})\|^{2}_{H_{t}^{-1}}$
$\leq\sum_{t\in\widetilde{{\mathcal{T}}}^{w}_{T+1}}\sum_{a\in S_{t}}\|\phi(x_{t},a)-\phi(x_{t},\bar{a}_{t})\|^{2}_{H_{t}^{-1}}$ ($\tilde{a}_{t},\bar{a}_{t}\in S_{t}$ by Eqn. (9))
$\leq\frac{2K^{2}}{\kappa}\sum_{t\in\widetilde{{\mathcal{T}}}^{w}_{T+1}}\sum_{a\in S_{t}}\|\phi(x_{t},a)-\phi(x_{t},\bar{a}_{t})\|^{2}_{\big(\Lambda^{w}_{t}\big)^{-1}}$ (Eqn. (D.14))
$\leq\frac{2K^{2}}{\kappa}\sum_{t=1}^{T}\min\left\{1,\sum_{a\in S_{t}}\|\phi(x_{t},a)-\phi(x_{t},\bar{a}_{t})\|^{2}_{\big(\Lambda^{w}_{t}\big)^{-1}}\right\}$ ($t\notin{\mathcal{T}}_{0}$ and $\widetilde{{\mathcal{T}}}^{w}_{T+1}\subseteq[T]$)

### v2

@@ C.2 Proofs of Lemmas for Theorem 1
$\sum_{j=1}^{|S_{s}|}\nabla^{2}\ell^{(j)}_{s}(\widehat{\bm{\theta}}_{s})$ $=\sum_{j=1}^{|S_{s}|}\sum_{k=j}^{|S_{s}|}\sum_{k^{\prime}=j}^{|S_{s}|}\frac{\exp\left(\left(\phi(x_{s},\sigma_{sk})+\phi(x_{s},\sigma_{sk^{\prime}})\right)^{\top}\widehat{\bm{\theta}}_{s}\right)}{2\left(\sum_{k^{\prime}=j}^{|S_{s}|}\exp\left(\phi(x_{s},\sigma_{sk^{\prime}})^{\top}\widehat{\bm{\theta}}_{s}\right)\right)^{2}}\cdot z_{skk^{\prime}}z_{skk^{\prime}}^{\top}$
$=\frac{1}{2}\sum_{j=1}^{|S_{s}|}\sum_{k=j}^{|S_{s}|}\sum_{k^{\prime}=j}^{|S_{s}|}P_{s}(\sigma_{sk}|S_{s}^{(j)};\widehat{\bm{\theta}}_{s})P_{s}(\sigma_{sk^{\prime}}|S_{s}^{(j)};\widehat{\bm{\theta}}_{s})z_{skk^{\prime}}z_{skk^{\prime}}^{\top}$
$=\frac{1}{2}\sum_{j=1}^{|S_{s}|}\sum_{k=j}^{|S_{s}|}\sum_{k^{\prime}=j}^{|S_{s}|}\mathbb{E}_{(a,a^{\prime})\sim\widehat{P}_{s}^{(j)}\times\widehat{P}_{s}^{(j)}}\left[\big(\phi(x_{s},a)-\phi(x_{s},a^{\prime})\big)\big(\phi(x_{s},a)-\phi(x_{s},a^{\prime})\big)^{\top}\right]$
$=\sum_{j=1}^{|S_{s}|}\mathbb{E}_{\widehat{P}_{s}^{(j)}}\left[\big(\phi(x_{s},a)-\mathbb{E}_{\widehat{P}_{s}^{(j)}}[\phi(x_{s},a^{\prime})]\big)\big(\phi(x_{s},a)-\mathbb{E}_{\widehat{P}_{s}^{(j)}}[\phi(x_{s},a^{\prime})]\big)^{\top}\right]\succeq 0,$ (C.11)
@@
where $z_{skk^{\prime}}=\phi(x_{s},\sigma_{sk})-\phi(x_{s},\sigma_{sk^{\prime}})$. Therefore, $\sum_{j=1}^{|S_{s}|}\nabla^{2}\ell^{(j)}_{s}(\widehat{\bm{\theta}}_{s})$ is a random PSD matrix conditional on the assortment $S_{s}$, where the randomness arises from the realized ranking $\sigma_{s}$. Note that the true PL distribution $\mathbb{P}_{s}(\cdot|S_{s};\bm{\theta}^{\star})$ is measurable with respect to the filtration $\mathcal{F}_{s-1}=\bm{\sigma}(x_{1},S_{1},\bm{\sigma}_{1},\dots,x_{s},S_{s})$. Moreover, the maximum eigenvalue of the Hessian satisfies $\lambda_{\max}\left(\sum_{j=1}^{|S_{s}|}\nabla^{2}\ell^{(j)}_{s}(\widehat{\bm{\theta}}_{s})\right)\leq K$. Then, by the PSD matrices concentration lemma (Lemma E.3), with probability $1-\delta$, we have
@@
$H_{t}$ $\succeq\frac{1}{e^{2}}\sum_{s\in[t-1]\setminus{\mathcal{T}}^{w}}\sum_{j=1}^{|S_{s}|}\nabla^{2}\ell^{(j)}_{s}(\widehat{\bm{\theta}}_{s})+\lambda\mathbf{I}_{d}$
@@
$=\frac{K}{e^{2}}\sum_{s\in[t-1]\setminus{\mathcal{T}}^{w}}\frac{1}{K}\sum_{j=1}^{|S_{s}|}\nabla^{2}\ell^{(j)}_{s}(\widehat{\bm{\theta}}_{s})+\lambda\mathbf{I}_{d}$
$\succeq\frac{K}{3e^{2}}\left(\sum_{s\in[t-1]\setminus{\mathcal{T}}^{w}}\mathbb{E}_{\sigma\sim\mathbb{P}_{s}(\cdot|S_{s},\bm{\theta}^{\star})}\left[\frac{1}{K}\sum_{j=1}^{|S_{s}|}\nabla^{2}\ell^{(j)}_{s,\sigma}(\widehat{\bm{\theta}}_{s})\right]+\lambda\mathbf{I}_{d}\right)$ (Lemma E.3, with prob. $1-\delta$)
$=\frac{1}{3e^{2}}\left(\sum_{s\in[t-1]\setminus{\mathcal{T}}^{w}}\sum_{j=1}^{|S_{s}|}\mathbb{E}_{\sigma\sim\mathbb{P}_{s}(\cdot|S_{s},\bm{\theta}^{\star})}\left[\nabla^{2}\ell^{(j)}_{s,\sigma}(\widehat{\bm{\theta}}_{s})\right]+\lambda\mathbf{I}_{d}\right)=\frac{1}{3e^{2}}M_{t}.$
This concludes the proof of Lemma C.2. ∎
C.2.2 Proof of Lemma C.3
Proof of Lemma C.3.
By the definition of $M_{t}$ (C.2) and Equation (C.11), for $t\notin{\mathcal{T}}^{w}$, we have
$\operatorname{det}\left(M_{t+1}\right)$ $=\operatorname{det}\left(M_{t}+\sum_{j=1}^{|S_{t}|}\mathbb{E}_{\sigma\sim\mathbb{P}_{t}(\cdot|S_{t},\bm{\theta}^{\star})}\left[\nabla^{2}\ell^{(j)}_{t,\sigma}(\widehat{\bm{\theta}}_{t})\right]\right)$
$=\operatorname{det}\left(M_{t}+\sum_{j=1}^{|S_{t}|}\mathbb{E}_{\begin{subarray}{c}\sigma\sim\mathbb{P}_{t}(\cdot|S_{t},\bm{\theta}^{\star})\\ a\sim P_{t}(\cdot|S^{(j)}_{t,\sigma};\widehat{\bm{\theta}}_{t})\end{subarray}}\left[\bar{\phi}^{(j)}_{t}(a)\bar{\phi}^{(j)}_{t}(a)^{\top}\right]\right)$ (Eqn. (C.11) )
$=\operatorname{det}\left(M_{t}\right)\left(1+\sum_{j=1}^{|S_{t}|}\mathbb{E}_{\begin{subarray}{c}\sigma\sim\mathbb{P}_{t}(\cdot|S_{t},\bm{\theta}^{\star})\\ a\sim P_{t}(\cdot|S^{(j)}_{t,\sigma};\widehat{\bm{\theta}}_{t})\end{subarray}}\left[\|\bar{\phi}^{(j)}_{t}(a)\|_{M_{t}^{-1}}^{2}\right]\right)$
$\geq\operatorname{det}\left(\lambda\mathbf{I}_{d}\right)\prod_{s\in[t]\setminus{\mathcal{T}}^{w}}\left(1+\min\left\{1,\sum_{j=1}^{|S_{s}|}\mathbb{E}_{\begin{subarray}{c}\sigma\sim\mathbb{P}_{s}(\cdot|S_{s},\bm{\theta}^{\star})\\ a\sim P_{s}(\cdot|S^{(j)}_{s,\sigma};\widehat{\bm{\theta}}_{s})\end{subarray}}\left[\|\bar{\phi}^{(j)}_{s}(a)\|_{M_{s}^{-1}}^{2}\right]\right\}\right),$
where, for simplicity, we denote $\bar{\phi}^{(j)}_{t}(a)=\phi(x_{t},a)-\mathbb{E}_{a^{\prime}\sim P_{t}(\cdot|S^{(j)}_{t,\sigma};\widehat{\bm{\theta}}_{t})}[\phi(x_{t},a^{\prime})]$.
For $t\in{\mathcal{T}}^{w}$, it is clear that $\operatorname{det}\left(M_{t+1}\right)=\operatorname{det}\left(M_{t}\right)$.
Hence, using the fact that $a\leq 2\log(1+a)$ for any $a\in[0,1]$, we get
$\sum_{t\notin{\mathcal{T}}^{w}}$ $\min\left\{1,\sum_{j=1}^{|S_{t}|}\mathbb{E}_{\begin{subarray}{c}\sigma\sim\mathbb{P}_{t}(\cdot|S_{t},\bm{\theta}^{\star})\\ a\sim P_{t}(\cdot|S^{(j)}_{t,\sigma};\widehat{\bm{\theta}}_{t})\end{subarray}}\left[\|\bar{\phi}^{(j)}_{t}(a)\|_{M_{t}^{-1}}^{2}\right]\right\}$
$\leq 2\sum_{t\notin{\mathcal{T}}^{w}}\log\left(1+\min\left\{1,\sum_{j=1}^{|S_{t}|}\mathbb{E}_{\begin{subarray}{c}\sigma\sim\mathbb{P}_{t}(\cdot|S_{t},\bm{\theta}^{\star})\\ a\sim P_{t}(\cdot|S^{(j)}_{t,\sigma};\widehat{\bm{\theta}}_{t})\end{subarray}}\left[\|\bar{\phi}^{(j)}_{t}(a)\|_{M_{t}^{-1}}^{2}\right]\right\}\right)$
$\leq 2\log\left(\frac{\operatorname{det}\left(M_{T+1}\right)}{\operatorname{det}\left(\lambda\mathbf{I}_{d}\right)}\right)$
$\leq 2d\log\left(1+\frac{4KT}{d\lambda}\right),$
@@
$\operatorname{det}\left(M_{T+1}\right)$ $\leq\left(\frac{\lambda_{1}+\cdots+\lambda_{d}}{d}\right)^{d}$ ($\lambda_{1},\cdots,\lambda_{d}$ are eigenvalues of $M_{T+1}$, AM-GM ineq.)
$\leq\left(\frac{\operatorname{trace}(M_{T+1})}{d}\right)^{d}$
$=\left(\frac{\lambda d+\sum_{t=1}^{T}\sum_{j=1}^{|S_{t}|}\mathbb{E}\|\bar{\phi}^{(j)}_{t}(a)\|_{2}^{2}}{d}\right)^{d}\leq\left(\lambda+\frac{KT}{d}\right)^{d}.$ ($\mathbb{E}\|\bar{\phi}^{(j)}_{t}(a)\|_{2}^{2}\leq 1$)
This concludes the proof of Lemma C.3. ∎
C.2.3 Proof of Lemma C.4
Proof of Lemma C.4.
Let
$(\tilde{a}_{t},\bar{a}_{t})=\mathop{\mathrm{argmax}}_{a,a^{\prime}\in\mathcal{A}}\dot{\mu}\left(\big(\phi(x_{t},a)-\phi(x_{t},a^{\prime})\big)^{\top}\widehat{\bm{\theta}}_{t}\right)\|\phi(x_{t},a)-\phi(x_{t},a^{\prime})\|_{H_{t}^{-1}}^{2},$
where $\mu(w)=\frac{1}{1+e^{-w}}$ denote the sigmoid function. Note that, by our assortment selection rule in Equation (11), the two actions $a,a^{\prime}\in\mathcal{A}$ that maximize $\dot{\mu}\left(\big(\phi(x_{t},a)-\phi(x_{t},a^{\prime})\big)^{\top}\widehat{\bm{\theta}}_{t}\right)\|\phi(x_{t},a)-\phi(x_{t},a^{\prime})\|^{2}_{H_{t}^{-1}}$ are always included in $S_{t}$. It is because
$\dot{\mu}$ $\left(\big(\phi(x_{t},a)-\phi(x_{t},a^{\prime})\big)^{\top}\widehat{\bm{\theta}}_{t}\right)\|\phi(x_{t},a)-\phi(x_{t},a^{\prime})\|^{2}_{H_{t}^{-1}}$
$=2\mathbb{E}_{a\sim P_{t}(\cdot|\{a,a^{\prime}\};\widehat{\bm{\theta}}_{t})}\left[\left\|\phi(x_{t},a)-\mathbb{E}_{a^{\prime}\sim P_{t}(\cdot|\{a,a^{\prime}\};\widehat{\bm{\theta}}_{t})}[\phi(x_{t},a^{\prime})]\right\|_{H_{t}^{-1}}^{2}\right]$
$=2\mathbb{E}_{\begin{subarray}{c}\sigma\sim\mathbb{P}_{t}(\cdot|\{a,a^{\prime}\};\widehat{\bm{\theta}}_{t})\\ a\sim P_{t}(\cdot|\{a,a^{\prime}\};\widehat{\bm{\theta}}_{t})\end{subarray}}\left[\left\|\phi(x_{t},a)-\mathbb{E}_{a^{\prime}\sim P_{t}(\cdot|\{a,a^{\prime}\};\widehat{\bm{\theta}}_{t})}[\phi(x_{t},a^{\prime})]\right\|_{H_{t}^{-1}}^{2}\right]=2f_{t}(\{a,a^{\prime}\}),$
which directly implies that $\tilde{a}_{t},\bar{a}_{t}\in S_{t}$.
For simplicity, we denote $z_{skk^{\prime}}=\phi(x_{s},\sigma_{sk})-\phi(x_{s},\sigma_{sk^{\prime}})$. Then, by the definition of $H_{t}$, we have
@@
$\succeq\frac{e^{-4B}}{2K^{2}}\sum_{s=1}^{t-1}\sum_{j=1}^{|S_{s}|}\sum_{k=j}^{|S_{s}|}\sum_{k^{\prime}=j}^{|S_{s}|}z_{skk^{\prime}}z_{skk^{\prime}}^{\top}+\lambda\mathbf{I}_{d}$
$\succeq\frac{e^{-4B}}{2K^{2}}\sum_{s=1}^{t-1}\big(\phi(x_{s},\tilde{a}_{s})-\phi(x_{s},\bar{a}_{s})\big)\big(\phi(x_{s},\tilde{a}_{s})-\phi(x_{s},\bar{a}_{s})\big)^{\top}+\lambda\mathbf{I}_{d}$ ( $\tilde{a}_{s},\bar{a}_{s}\in S_{s}$)
$\succeq\frac{e^{-4B}}{2K^{2}}\Bigg(\sum_{s\in{\mathcal{T}}^{w}_{t}}\dot{\mu}\!\left(\big(\phi(x_{s},\tilde{a}_{s})-\phi(x_{s},\bar{a}_{s})\big)^{\top}\widehat{\bm{\theta}}_{s}\right)\big(\phi(x_{s},\tilde{a}_{s})-\phi(x_{s},\bar{a}_{s})\big)\big(\phi(x_{s},\tilde{a}_{s})-\phi(x_{s},\bar{a}_{s})\big)^{\top}$
$\quad\quad\quad\quad+\lambda\mathbf{I}_{d}\Bigg),$ (C.12)
where the last inequality holds since $\dot{\mu}\!\left(\big(\phi(x_{s},a)-\phi(x_{s},a^{\prime})\big)^{\top}\widehat{\bm{\theta}}_{s}\right)\leq\frac{1}{4}$. In addition, we denote ${\mathcal{T}}^{w}_{t}={\mathcal{T}}^{w}\setminus\{t,\dots,T\}$ in the last inequality. To better presentation, we define
$\Lambda^{w}_{t}:=$ $\sum_{s\in{\mathcal{T}}^{w}_{t}}\dot{\mu}\!\left(\big(\phi(x_{s},\tilde{a}_{s})-\phi(x_{s},\bar{a}_{s})\big)^{\top}\widehat{\bm{\theta}}_{s}\right)\big(\phi(x_{s},\tilde{a}_{s})-\phi(x_{s},\bar{a}_{s})\big)\big(\phi(x_{s},\tilde{a}_{s})-\phi(x_{s},\bar{a}_{s})\big)^{\top}$
$+\lambda\mathbf{I}_{d}.$
On the other hand, by setting $\lambda\geq 1$, for any $t$, we can ensure that